OpenStudy (anonymous):

Given A(3, 4), B(4, 5), and C(3, 5), the length of AB + BC + AC is greater than:Answer 6 5 4 3

OpenStudy (dumbcow):

AB = sqrt(1^2 +1^2) = sqrt(2) BC = sqrt(1^2 +0^2) = 1 AC = sqrt(0^2 +1^2) = 1 1+1+sqrt(2) = 3.41 > 3

OpenStudy (anonymous):

yep